# Multilevel neural simulation-based inference

### v2

@@ B.1 The g-and-k distribution
where $\Phi^{-1}$ is the quantile function of the standard normal distribution and $\text{erf}^{-1}$ is the inverse of error function. We set the prior distribution to be a tensor product of marginal distributions on each parameter: $\theta_{1},\theta_{2},\theta_{3}\sim\text{Unif}([0,3]^{3})$ and $\theta_{4}\sim\text{Unif}([0,\exp(0.5)])$. Note that we always resort to an approximation method for the evaluation of $\text{erf}^{-1}(\cdot)$. For the high-fidelity simulator, we use an accurate approximation implemented in Scipy [Virtanen et al., 2020]. For the low-fidelity simulator, we use a Taylor expansion of $\text{erf}^{-1}$ up to the third order:
@@
An advantage of the g-and-k distribution is that its density can be approximated numerically almost exactly. Following Rayner and MacGillivray [2002], we first numerically obtain $G_{\theta}^{-1}(x)$ by solving numerically for $x_{i}-G_{\theta}(u_{i})=0$ using a root-solving algorithm [Press, 2007]111 Note that since $G_{\theta}(u)$ is a quantile function, $F_{\theta}(x):=G_{\theta}^{-1}(x)$ is a cumulative distribution function., then we obtain the density function by taking $\hat{p}(x\,|\,\theta):=F^{\prime}_{\theta}(x)=\partial F_{\theta}(x)/\partial x$. For this second step, we typically use a finite difference approximation. Seed-matching is simply done by using same the $u$ and $\theta$.
@@
We use NSF with $3$ bins, span of [3,3][-3,3] and $3$ coupling layers. The conditioner for the NSF is a MLP with $2$ hidden layers of $50$ units, and $10\%$ dropout, trained for $800$ epochs. For each true parameter values, we produce $m=1000$ i.i.d samples and obtain quantile based $4$-dimensional summary statistics following Prangle [2016].
@@ B.3 Ornstein–Uhlenbeck process
We use NSF with $2$ bins, span of [2,2][-2,2] and $2$ coupling layers. The conditioner for the NSF is a MLP with $1$ hidden layer of $20$ units, and $20\%$ dropout. For ML-NPE, we trained the network for $500$ epochs. For TL-NPE, we set a high maximum number of epochs and used on early stopping. The stopping criterion is based on the validation loss computed on $10\%$ of the data, with a patience parameter controlling how many epochs to wait before stopping. For each true parameter values, we produce $m=1$ sample and use $5$ representative data points as a summary statistics. The subsamples are taken at equal interval in log space such that we take $0,3,10,31,99$th data points.
@@
We train each ML-NPE and TL-NPE $20$ times hence the boxplot is over $20$ data points. For each trained network, we compute the median NLPD across $500$ simulated datasets. The use of the median, rather than the mean, is motivated by the presence of occasional extreme NLPD values in a few simulations. To estimate the KLD, we draw $2000$ samples from the reference density, and evaluate the densities at each sampled value of $\theta$ for the both target and reference NPE, and compute the KLD for each simulation. The final KLD value for each network is obtained by averaging the results over $500$ simulations.
@@
Figure 7: Additional results for the OUP experiment. (a) Example of one seed-matched sample from high and low fidelity simulators. (b) Recovery plot for the reference NPE posterior with $n=10^{4}$. (c)-(f) Recovery plots for ML-NPE and TL-NPE for $n_{1}=10$ and $n_{1}=100$. For TL-NLE, we picked the result corresponding to the best performing hyperparameter of their stopping criterion (i.e. patience =1=1 for $n_{1}$ = 10 and patience =20=20 for $n_{1}=100$) to visualise. For $n_{1}=10$, ML-NPE achieves better performance on $\gamma$ and $\mu$. For $n_{1}=100$, it achieves superior performance on $\gamma$ and competitive results on the remaining two. Recall that, unlike TL-NPE, ML-NPE does not have any hyperparameter to tune.
@@ B.3.1 Experiment with different dimension of $\theta$
We now conduct an experiment to explore a failure mode of our method. We include an additional parameter in the high fidelity simulator, termed “initial value” $\theta_{4}=x_{0}\sim\text{uniform}([0,4])$, instead of fixing it at $x_{0}=2.0$ as we did previously. For this experiment, we picked $n_{0}=1,000$ and $n_{1}=100$. We additionally trained the conditional density estimator using only $n=100$ data from the high fidelity simulator. All the other settings remain the same.
@@
ML-NPE TL-NPE NPE (high only)
NLPD $\downarrow$ -0.18 -0.89 -1.21
(0.04) (0.52) (0.53)
KLD $\downarrow$ 0.98 1.24 1.05
(0.08) (0.40) (0.14)
We observe that although our method shows a slight improvement in KLD, it notably underperforms in NLPD compared to using only high-fidelity data. This may be due to our training objective being more unstable and therefore more sensitive to large differences in simulator outputs introduced by the additional parameter $x_{0}$.
@@ B.4 Toggle switch model
We sample $m=500$ from both high fidelity simulator and each approximated densities, and calculate MMD over $5000$ simulations. The length scale is estimated by the median heuristic.
@@
Figure 8: Additional results for the Toggle-switch experiment. (a) Histograms of $1000$ seed-matched samples from the high-, the medium-, and the low-fidelity simulators. Note that the difference between the low- and the medium-fidelity simulator is larger than the difference between the medium- and the high-fidelity simulator. This suggests that we should take $n_{1}>n_{2}$. (b) Approximated density and samples from high-fidelity simulator across four simulations. ML-NLE demonstrates superior performance compared to the NLE baselines trained exclusively on high-, medium-, and low-fidelity data with same total simulation budget. In particular, the density estimated by ML-NLE aligns more closely with high-fidelity simulations, highlighting the improved accuracy of our method.
@@ B.5 Cosmological simulations
We use the CAMELS simulation suite [Villaescusa-Navarro et al., 2021, 2023], a benchmark dataset for machine learning in astrophysics, to study multi-fidelity simulation-based inference in cosmology. The simulation is one of the most expensive cosmological suites ever run; a small fraction of the next generation are being run on the UK DIRAC HPC Facility with 15M CPUh.
@@
In this setup, low-fidelity simulations are governed solely by $\theta$ and the initial condition seed $u$, while high-fidelity simulations additionally incorporate complex astrophysical processes (e.g., feedback), modelled via four extra parameters. These high-fidelity simulations are significantly more computationally expensive—often more than 100× slower to generate than low-fidelity ones. Given these limitations, cosmological analyses often rely on conservative data cuts to exclude small-scale modes (e.g., $k\gtrsim 0.1,h/\text{Mpc}$), where simulation inaccuracies are most pronounced [e.g. Jeffrey et al., 2025, Gatti et al., 2025]. Multi-fidelity approaches aim to mitigate such constraints by leveraging both inexpensive and high-fidelity simulations to improve the accuracy of cosmological inference.
Note that the idea of combining low- and high- fidelity simulations has previously been proposed in cosmology; see for example the work of Chartier et al. [2021], Chartier and Wandelt [2022] who use low-fidelity simulations to construct approximations of quantities of interest which can be used as control variates. This work differs from our proposed approach in that it is based on regression and approximations of quantities such as means and covariances of interest, rather than of the training objective of neural SBI.
@@
Figure 9: Additional results for the cosmological simulator experiment. (a) 1 seed-matched sample from high and low fidelity simulators. The one with noise is used to train the neural networks. (b)-(c) Recovery plot of ML-NPE and NPE. Adding low fidelity samples lead better recovery of the parameter.

### v4

@@ A.2 Proof of Theorem˜2
A.3 Extension of Theorem˜1 to the gradient
We provide an upper bound on the variance for each element of the gradient $\nabla_{\phi}\ell_{\text{MLMC}}(\phi)$, that is, on each partial derivative $\nabla_{\phi_{j}}\ell_{\text{MLMC}}(\phi)$ for $j\in\{1,\ldots,d_{\phi}\}$. The partial derivatives are given by
$\nabla_{\phi_{j}}\ell_{\text{MLMC}}(\phi)$ $=\nabla_{\phi_{j}}h_{0}(\phi)+\sum_{l=1}^{L}\nabla_{\phi_{j}}h_{l}(\phi)$
$=\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}\nabla_{\phi_{j}}f_{\phi}^{0}(u_{i}^{0},\theta_{i}^{0})+\sum_{l=1}^{L}\frac{1}{n_{l}}\sum_{i=1}^{n_{l}}(\nabla_{\phi_{j}}f^{l}_{\phi}(u_{i}^{l},\theta_{i}^{l})-\nabla_{\phi_{j}}f^{l-1}(u_{i}^{l},\theta_{i}^{l}))$
Theorem 3.
Let $\phi\in\Phi\subseteq\mathbb{R}^{d_{\Phi}}$ and suppose the following assumption hold in addition to A1-A2 in theorem˜1:
  1. (A3’)
    $\nabla_{\phi_{j}}\log\tilde{q}_{\phi}$ is continuously differentiable, locally $K^{j}_{\text{Lip}}(\phi)$-smooth and satisfies the growth condition $\|\nabla_{\phi_{j}}\nabla\log\tilde{q}_{\phi}(x_{1:m},\theta)\|_{2}\leq K^{j}_{\text{grow}}(\phi)(\sum_{i=1}^{m}\|x_{i}\|_{2}+\|\theta\|_{2}+1)$ for some $K^{j}_{\text{Lip}}(\phi),K^{j}_{\text{grow}}(\phi)>0$ for all $j=1,\dots,d_{\Phi}$.
Then, for $l\in\{1,\ldots,L\}$, $K_{0}^{j}(\phi),\ldots,K_{L}^{j}(\phi)>0$, $j\in\{1,\ldots,d_{\phi}\}$ independent of $n_{0},\ldots,n_{L}$, we have that:
$\text{Var}\left[\nabla_{\phi_{j}}h_{0}(\phi)\right]$ $\leq\frac{K^{j}_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U})}^{4}+1\right),$
$\text{and}\;\text{Var}\left[\nabla_{\phi_{j}}h_{l}(\phi)\right]$ $\leq\frac{K^{j}_{l}(\phi)}{n_{l}}\left(\ \|G^{l}-G^{l-1}\|_{W^{1,4}(\pi\times\mathbb{U})}^{2}\right)$
Proof.
The variance of each term can be expressed as
$\text{Var}\left[\nabla_{\phi_{j}}h_{0}(\phi)\right]$ $=\frac{1}{n_{0}}\text{Var}[\nabla_{\phi_{j}}f_{\phi}^{0}(z)]$ (48)
$\text{Var}\left[\nabla_{\phi_{j}}h_{l}(\phi)\right]$ $=\frac{1}{n_{l}}\text{Var}[\nabla_{\phi_{j}}f_{\phi}^{l}(z)-\nabla_{\phi_{j}}f_{\phi}^{l-1}(z)]$ (49)
Assume that $\nabla_{\phi_{j}}f_{\phi}$ is sufficiently regular, applying Poincaré inequality to ˜48 and ˜49 (as in ˜8 and ˜9) gives:
$\text{Var}[\nabla_{\phi_{j}}f_{\phi}^{0}(z)]$ $\leq K_{\text{Poin}}\mathbb{E}\left[\left\|\nabla_{z}\nabla_{\phi_{j}}f_{\phi}^{0}(z)\right\|_{2}^{2}\right]$ (50)
$\text{Var}[\nabla_{\phi_{j}}f_{\phi}^{l}(z)-\nabla_{\phi_{j}}f_{\phi}^{l-1}(z)]$ $\leq K_{\text{Poin}}\mathbb{E}\left[\left\|\nabla_{z}\nabla_{\phi_{j}}f_{\phi}^{l}(z)-\nabla_{z}\nabla_{\phi_{j}}f_{\phi}^{l-1}(z)\right\|_{2}^{2}\right]$ (51)
where the expectation is over $z$. We can simplify ˜50 as
$\mathbb{E}\left[\left\|\nabla_{z}\nabla_{\phi_{j}}f_{\phi}^{0}(z)\right\|_{2}^{2}\right]$ $=\mathbb{E}\left[\left\|\nabla_{z}\nabla_{\phi_{j}}\log\tilde{q}_{\phi}(g^{0}(z))\right\|_{2}^{2}\right]$ (52)
$=\mathbb{E}\left[\left\|\nabla_{z}g^{0}(z)\nabla\nabla_{\phi_{j}}\log\tilde{q}_{\phi}(g^{0}(z))\right\|_{2}^{2}\right]$ (53)
$\leq\mathbb{E}\left[\left\|\nabla_{z}g^{0}(z)\right\|_{2}^{4}\right]^{\frac{1}{2}}\mathbb{E}\left[\left\|\nabla\nabla_{\phi_{j}}\log\tilde{q}_{\phi}(g^{0}(z))\right\|_{2}^{4}\right]^{\frac{1}{2}}$ (54)
Here ˜53 is due to the chain rule and ˜54 follows ˜11 - ˜12. The bound for the first expectation is given by ˜20. For the second expectation, we have:
$\mathbb{E}\left[\left\|\nabla\nabla_{\phi_{j}}\log\tilde{q}_{\phi}(g^{0}(z))\right\|_{2}^{4}\right]$ $\leq K^{j}_{\text{score}}(\phi)\left(\left\|G^{0}\right\|_{W^{1,4}}^{4}+1\right)$ (55)
by Assumption (A3’) and following ˜21 - ˜24. Combining ˜48, ˜50, ˜54, ˜20, and ˜55 gives:
$\text{Var}[\nabla_{\phi_{j}}h_{0}(\phi)]\leq\frac{K_{0}^{j}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|_{W^{1,4}}^{4}+1\right)$ (56)
where $K^{j}_{0}(\phi)$ is a constant depending on $\phi$ and $j$, independent of $n_{0}$. Now, we derive an upper bound on $\text{Var}\left[\nabla_{\phi_{j}}h_{l}(\phi)\right]$. Following ˜26 - ˜30, we have:
$\mathbb{E}\left[\left\|\nabla_{z}\nabla_{\phi_{j}}f_{\phi}^{l}(z)-\nabla_{z}\nabla_{\phi_{j}}f_{\phi}^{l-1}(z)\right\|_{2}^{2}\right]$ (57)
$\leq 2\Bigg[\left(\mathbb{E}\left[\left\|\nabla_{z}g^{l}(z)\right\|_{4}^{2}\right]\right)^{\frac{1}{2}}\left(\mathbb{E}\left[\left\|\nabla\nabla_{\phi_{j}}\log\tilde{q}_{\phi}(g^{l}(z))-\nabla\nabla_{\phi_{j}}\log\tilde{q}_{\phi}(g^{l-1}(z))\right\|_{2}^{4}\right]\right)^{\frac{1}{2}}$
$\quad+\mathbb{E}\left[\left\|\nabla_{z}g^{l}(z)-\nabla_{z}g^{l-1}(z)\right\|_{2}^{4}\right]^{\frac{1}{2}}\mathbb{E}\left[\left\|\nabla\nabla_{\phi_{j}}\log\tilde{q}_{\phi}(g^{l-1}(z))\right\|_{2}^{4}\right]^{\frac{1}{2}}\Bigg]$ (58)
Bound for the first expectation is given by ˜31 and the third expectation is given by ˜32. For the forth expectation, from ˜55 (replacing 0 with $l$) and Assumption (A2), we have:
$\mathbb{E}\left[\left\|\nabla\nabla_{\phi_{j}}\log\tilde{q}_{\phi}\left(g^{l-1}(z)\right)\right\|_{2}^{4}\right]\leq K^{j}_{\text{score}}(\phi)\left(S_{l-1}^{4}+1\right)$ (59)
For the second expectation, following ˜36-˜40, and using our local Lipschitz property, we get:
$\mathbb{E}\left[\left\|\nabla\nabla_{\phi_{j}}\log\tilde{q}_{\phi}(g^{l}(z))-\nabla\nabla_{\phi_{j}}\log\tilde{q}_{\phi}(g^{l-1}(z))\right\|_{2}^{4}\right]$ (60)
$\leq K^{j}_{\text{score-diff}}(\phi,\delta)\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}}^{4}$ (61)
where $K^{j}_{\text{score-diff}}(\phi)$ combines all the constant, independent of $n_{l}$ and the difference in the simulators.
Finally, applying the bounds for the each expectation; ˜31, ˜61, ˜32, and ˜59, to ˜58, and combining it with ˜49 gives:
$\text{Var}[\nabla_{\phi_{j}}h_{0}(\phi)]$ $\leq\frac{2K_{\text{Poin}}}{n_{l}}\big[K^{\frac{1}{2}}_{\text{grad}}(S_{l}^{4}+1)^{\frac{1}{2}}(K^{j}_{\text{score-diff}}(\phi))^{\frac{1}{2}}\|G^{l}-G^{l-1}\|_{W^{1,4}}^{2}$
$+K^{\frac{1}{2}}_{\text{grad}}\left(\|G^{l}-G^{l-1}\|_{W^{1,4}}^{4}\right)^{\frac{1}{2}}(K^{j}_{\text{score}}(\phi))^{\frac{1}{2}}(S^{4}_{l-1}+1)^{\frac{1}{2}}\big]$ (62)
$\leq\frac{K^{j}_{l}(\phi)}{n_{l}}\left(\ \|G^{l}-G^{l-1}\|_{W^{1,4}}^{2}\right)$ (63)
where $K_{l}$ combines all the constant. Combining ˜56 and ˜63 gives a bound on the variance of the partial derivative:
$\text{Var}[\nabla_{\phi_{j}}\ell_{\text{MLMC}}(\phi)]$ $\leq\frac{K^{j}_{0}(\phi)}{n_{0}}\left(\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U})}^{4}+1\right)+\sum_{l=1}^{L}\frac{K^{j}_{l}(\phi)}{n_{l}}\|G^{l}-G^{l-1}\|_{W^{1,4}(\pi\times\mathbb{U})}^{2}$ (64)
@@ B.1 The g-and-k distribution
where $\Phi^{-1}$ is the quantile function of the standard normal distribution and $\text{erf}^{-1}$ is the inverse of error function. We set the prior distribution to be a tensor product of marginal distributions on each parameter: $\theta_{1},\theta_{2},\theta_{3}\sim\text{Unif}([0,3]^{3})$ and $\theta_{4}\sim\text{Unif}([0,\exp(0.5)])$. Note that we always resort to an approximation method for the evaluation of $\text{erf}^{-1}(\cdot)$. For the high-fidelity simulator, we use an accurate approximation implemented in Scipy [Virtanen and others, 2020]. For the low-fidelity simulator, we use a Taylor expansion of $\text{erf}^{-1}$ up to the third order:
@@
An advantage of the g-and-k distribution is that its density can be approximated numerically almost exactly. Following Rayner and MacGillivray [2002], we first numerically obtain $F_{\theta}(x)=G_{\theta}^{-1}(x)$ by solving numerically for $x_{i}-G_{\theta}(u_{i})=0$ using a root-solving algorithm [Press, 2007] [^1] Note that since $G_{\theta}(u)$ is a quantile function, $F_{\theta}(x):=G_{\theta}^{-1}(x)$ is a cumulative distribution function., then we obtain the density function by taking $\hat{p}(x\,|\,\theta):=F^{\prime}_{\theta}(x)=\partial F_{\theta}(x)/\partial x$. For this second step, we typically use a finite difference approximation. Seed-matching is simply done by using same the $u$ and $\theta$.
@@
We use NSF with $3$ bins, span of [3,3][-3,3] and $3$ coupling layers. The conditioner for the NSF is a MLP with $2$ hidden layers of $50$ units, and $10\%$ dropout, trained for $800$ epochs. For each true parameter values, we produce $m=1000$ iid samples and obtain quantile based $4$-dimensional summary statistics following Prangle [2016].
@@ B.3 Ornstein–Uhlenbeck process
We use NSF with $2$ bins, span of [2,2][-2,2] and $2$ coupling layers. The conditioner for the NSF is a MLP with $1$ hidden layer of $32$ units, and $10\%$ dropout. For ML-NPE, we trained the network for $500$ epochs. For TL-NPE, we set a high maximum number of epochs and used on early stopping. The stopping criterion is based on the validation loss computed on $20\%$ of the data, with a patience parameter controlling how many epochs to wait before stopping. For each true parameter values, we produce $m=1$ sample and use $5$ representative data points as a summary statistics. The subsamples are taken at equal interval in log space such that we take $0,3,10,31,99$th data points.
@@
We train each ML-NPE and TL-NPE $20$ times. For each trained network, we compute the NLPD and KL divergence against reference posterior across $500$ simulated datasets.
@@
Figure 7: Additional results for the OUP experiment. (a) Example of one seed-matched sample from high and low-fidelity simulators. (b) Recovery plot for the reference NPE posterior with $n=10^{4}$. (c)-(f) Recovery plots for ML-NPE and TL-NPE for $n_{1}=10$ and $n_{1}=100$.
((a)) $n_{1}=10$
((b)) $n_{1}=100$
Figure 8: Figure˜8 with all the data, without removing outliers. Outliers are identified using IQR method with factor 44, which removes only extreme outliers.
@@ B.3.1 Experiment with different dimension of $\theta$
We now conduct an experiment to explore a failure mode of our method. We include an additional parameter in the high-fidelity simulator, termed “initial value” $\theta_{4}=x_{0}\sim\text{uniform}([0,4])$, instead of fixing it at $x_{0}=2.0$ as we did previously. For this experiment, we picked $n_{0}=1,000$ and $n_{1}=100$. We additionally trained the conditional density estimator using only $n=100$ data from the high-fidelity simulator. All the other settings remain the same.
@@
ML-NPE NPE (high only)
NLPD $\downarrow$ -0.18 -1.21
(0.04) (0.53)
We observe that it notably underperforms in NLPD compared to using only high-fidelity data. This may be due to our training objective being more unstable and therefore more sensitive to large differences in simulator outputs introduced by the additional parameter $x_{0}$.
@@ B.4 Toggle switch model
We sample $m=500$ from both high-fidelity simulator and each approximated densities, and calculate MMD over $5000$ simulations. The length scale is estimated by the median heuristic.
@@
Figure 9: Additional results for the Toggle-switch experiment. (a) Histograms of $1000$ seed-matched samples from the high-, the medium-, and the low-fidelity simulators. Note that the difference between the low- and the medium-fidelity simulator is larger than the difference between the medium- and the high-fidelity simulator. This suggests that we should take $n_{1}>n_{2}$. (b) Approximated density and samples from high-fidelity simulator across four simulations. ML-NLE demonstrates superior performance compared to the NLE baselines trained exclusively on high-, medium-, and low-fidelity data with the same total simulation budget. In particular, the density estimated by ML-NLE aligns more closely with high-fidelity simulations, highlighting the improved accuracy of our method.
@@ B.5 Cosmological simulations
We use the CAMELS simulation suite [Villaescusa-Navarro and others, 2021, 2023], a benchmark dataset for machine learning in astrophysics, to study multi-fidelity simulation-based inference in cosmology. The simulation is one of the most expensive cosmological suites ever run; a small fraction of the next generation are being run on the UK DIRAC HPC Facility with 15M CPUh.
@@
In this setup, low-fidelity simulations are governed solely by $\theta$ and the initial condition seed $u$, while high-fidelity simulations additionally incorporate complex astrophysical processes (e.g., feedback), modelled via four extra parameters. These high-fidelity simulations are significantly more computationally expensive—often more than 100× slower to generate than low-fidelity ones. Given these limitations, cosmological analyses often rely on conservative data cuts to exclude small-scale modes (e.g., $k\gtrsim 0.1,h/\text{Mpc}$), where simulation inaccuracies are most pronounced [e.g. Jeffrey and others, 2025, Gatti and others, 2025]. Multi-fidelity approaches aim to mitigate such constraints by leveraging both inexpensive and high-fidelity simulations to improve the accuracy of cosmological inference.
Note that the idea of combining low- and high- fidelity simulations has previously been proposed in cosmology; see for example the work of Chartier et al. [2021], Chartier and Wandelt [2022] who use low-fidelity simulations to construct approximations of quantities of interest which can be used as control variates. This work differs from our proposed approach in that it targets quantities such as means and covariances, rather than f the training objective of neural SBI.
@@
Figure 10: Additional results for the cosmological simulator experiment. (a) 1 seed-matched sample from high and low-fidelity simulators. The one with noise is used to train the neural networks. (b)-(c) Recovery plot of ML-NPE and NPE. Adding low-fidelity samples lead better recovery of the parameter.